OpenStudy (96mertcank96):

PLEASE HELP!!! 3/8 - 9/4 / [(- 5/4 + (- 1/10)] = ??? PEMDAS

OpenStudy (anonymous):

2.041666667

OpenStudy (anonymous):

i mean 1.388888889

OpenStudy (96mertcank96):

Whats ur work though?

OpenStudy (jonnyvonny):

Is it (3/8-(9/4))/((-5/4)+(-1/10))?

OpenStudy (96mertcank96):

yes

OpenStudy (jonnyvonny):

So, first you need to find common denominators for the fractions that are adding/subtracting each other; ((3/8)-(9/4)) has the common denominator 8.\[((3/8)-(9/4))\left| \right| (-9/4)*(2/2)\rightarrow (-18/8)\rightarrow ((3/8)-(18/8))=(-15/8)\] (-5/4) and (-1/10) have the common denominator is 40 (just multiply the denominators together). \[(-5/4)+(-1/10)\rightarrow (-5/4)*(10/10) + (-1/10)*(4/4)\rightarrow(-50/40)+(-4/40)\]\[=(-54/40)\] Now divide: \[(-15/8)/(-54/40)\rightarrow (-15/8)*(-40/58)\] Multiply straight across and that's your answer :)